OpenStudy (anonymous):

Let u = <7, -3>, v = <-9, 5>. Find 4u - 3v.

OpenStudy (jdoe0001):

that'd be just plain addition of vectors

OpenStudy (anonymous):

I didnt read about that :/

OpenStudy (anonymous):

8, -6?

OpenStudy (jdoe0001):

a * < b, c> => < a*b, a*c> # scalar multiplication <b , c > + <d, e> => <b+d, c+e> # addition

OpenStudy (anonymous):

I'm sorry I dont understnad

OpenStudy (jdoe0001):

4u => 4 < 7, -3 > # scalar multiplication

OpenStudy (anonymous):

ok...so gimme a sec please

OpenStudy (anonymous):

i got 1, 12...thats not right

OpenStudy (jdoe0001):

you have 4u - 3v so you do the "scalar multiplication" first then do the "vector addition" operation on the products

OpenStudy (anonymous):

...ill try again...

OpenStudy (anonymous):

i dont get it...4(7,3) -3(-9,5)

OpenStudy (anonymous):

is the answer then 55, -27

OpenStudy (jdoe0001):

\(\large \begin{matrix} a \times <b, c> &\implies & < a \times b, a\times c> &\textit{ scalar multiplication}\\ <b , c > + <d, e> & \implies & <b+d, c+e> &\textit{vector addition} \end{matrix}\)

OpenStudy (anonymous):

where did d and e come from :(

OpenStudy (jdoe0001):

hehe, it's just a notation for the operations

OpenStudy (anonymous):

:((((

OpenStudy (jdoe0001):

that's how you'd multiply a scalar that's how you'd add 2 vectors

OpenStudy (anonymous):

It doesnt make sense

OpenStudy (jdoe0001):

\( \begin{matrix} \color{red}{a} \times <b, c> &\implies & < \color{red}{a} \times b, \color{red}{a}\times c> &\textit{ scalar multiplication}\\ <\color{blue}{b} ,\color{blue}{ c} > + <\color{green}{d}, \color{green}{e}> & \implies & <\color{blue}{b}+\color{green}{d}, \color{blue}{c}+\color{green}{e}> &\textit{vector addition} \end{matrix}\) see it now?

OpenStudy (anonymous):

no

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=pNMrYACjHXQ

OpenStudy (anonymous):

so i just need the botton formula you keep giving me

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so the answer is 55, -27

OpenStudy (anonymous):

right ?

OpenStudy (jdoe0001):

hmm, I got something else, the -27 is correct though

OpenStudy (anonymous):

that is the only answer possible...

OpenStudy (anonymous):

28+27 = 55

OpenStudy (anonymous):

Wanna help with another?

OpenStudy (jdoe0001):

28 + (-27) => 1

OpenStudy (anonymous):

please?

OpenStudy (anonymous):

but its positive 27 becaise -9 and -3

OpenStudy (jdoe0001):

ohhh nevermind, yes 55, I see what happened

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so wanna help again?

OpenStudy (jdoe0001):

I can try

OpenStudy (anonymous):

ok :) Let u = <-3, -5>, v = <-3, 1>. Find u + v

OpenStudy (anonymous):

wait!

OpenStudy (anonymous):

-6, -4 ???

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

THANK YOUUUU

OpenStudy (jdoe0001):

yw